# RiemannFormer: A Framework for Attention in Curved Spaces

### v2

@@ Abstract
This research endeavors to unlock the further potential of transformer-based architectures by providing a geometric interpretation of the self-attention mechanism. In our framework, we reframe attention as interactions on a potential Attention Manifold. Queries and keys are treated as vectors within the manifold’s intrinsic geometry, residing in the tangent bundle, where the geometry of each tangent space is governed by a Riemannian metric. This metric is the fundamental object that defines both the local inner product for similarity computation and the connection that determines parallel transport. In stark contrast, we propose that value vectors are sections of a separate, trivial vector bundle, decoupling the geometric alignment of $Q/K$ from the feature aggregation of $V$. Our primary contribution is a novel, first-principles construction of this parallel transport using Lie group theory, where for 2D data, we derive a positional encoding from the exceptional isomorphism $\mathfrak{so}(4)\cong\mathfrak{su}(2)\oplus\mathfrak{su}(2)$. This allows us to factorize 4D rotations into two commuting isoclinic rotations aligned with the spatial dimensions. This dual-bundle structure provides a principled foundation for our explicit mechanism that highlights a neighborhood by attenuating remote values. Experimental results demonstrate that our modules deliver significant performance improvements over strong baselines, and further evaluations on visual and large language models are underway.
@@ 1 Introduction
The transformer block operates on a sequence of tokens, each represented as a vector of fixed dimensionality $d_{\text{model}}$. For a sequence of $L$ tokens, the input to the block is a matrix $X\in\mathbb{R}^{L\times d_{\text{model}}}$, where each row corresponds to a token’s embedding. The transformer block processes this matrix through its core component, Multi-Head Self-Attention (MHSA). The MHSA mechanism is the cornerstone of the transformer, enabling it to aggregate information globally across all tokens in the input sequence. Unlike convolutional neural networks (CNNs), which rely on local receptive fields, or recurrent neural networks (RNNs), which process sequences sequentially, self-attention allows each token to attend to all other tokens simultaneously, capturing long-range dependencies efficiently. Specifically, the attention mechanism computes a score for each pair of tokens by taking the dot product of their query and key vectors, scaled to prevent large values that could destabilize training: $\text{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d_{k}}}\right)\mathbf{V}$. The dot product $\mathbf{QK}^{\top}\in\mathbb{R}^{L\times L}$ measures the similarity between the query vector of one token and the key vectors of all tokens, producing an attention score matrix. The softmax function normalizes the scores across the key dimension, converting them into weights that sum to $1$ for each query. This ensures that the attention mechanism focuses on the most relevant tokens. The resulting weights are used to compute a weighted sum of the value vectors, producing a new representation for each token that incorporates information from all other tokens. The output of the attention mechanism is a matrix, where each row is a context-aware representation of the corresponding input token. Moreover, the attention mechanism is permutation-equivariant, meaning the output does not depend on the order of the input tokens unless positional information is explicitly added.
@@ 2 Methodology
Consider a sequence comprising $L$ input tokens, with their embeddings represented as $\{\mathbf{x}_{i}\}_{i=1}^{L}$, where $\mathbf{x}_{i}\in\mathbb{R}^{D}$ denotes the $D$-dimensional embedding vector for the $i$-th token, devoid of positional information. In this paper, we adopt an alternative perspective: we posit the existence of a curved space $\mathcal{M}$ wherein the encoding of a token sequence corresponds to identifying a sequence of points $\{\mathbf{p}_{i}\}$, or a grid of points $\{\mathbf{p}_{ij}\}$ for image data, etc.
@@ 2.2 Parallel Transport
The above relationship of these matrices is what we must consider when predefining them. The Riemannian metric $\mathbf{M}_{i}$ is a symmetric positive matrix. To facilitate the model to effectively learn the linear transformations $\mathbf{T}_{i}$, we first decompose them into two components, including scaling and length-preserving in $\mathbb{R}^{D}$: $\mathbf{T}_{i}=\mathbf{S}_{i}\mathbf{R}_{i}$. The introduction of scaling part $\mathbf{S}_{i}$ is motivated by the characteristics of the curved space and the existence of Riemannian metric. The scaling part is represented by a symmetric positive matrix, and its simple form is a diagonal matrix or a scalar matrix $s\mathbf{I}$, where $\mathbf{I}$ is the identity matrix. For instance, we set $\mathbf{S}_{i}=\mathbf{M}_{i}^{-\frac{1}{2}}$, then we have $\mathbf{T}_{i}=\mathbf{M}_{i}^{-\frac{1}{2}}\mathbf{R}_{i}$. $\mathbf{R}_{i}$ is an orthogonal matrix used for length-preserving, and a rotation is naturally a candidate. A rotation in $n$-dimensional Euclidean space is a linear transformation that preserves lengths of vectors, and it can be represented by an $D\times D$ orthogonal matrix in the special orthogonal group $\mathbf{SO}(D)$.
@@
where $\langle\partial_{i},\partial_{i}\rangle_{\mathbf{p}_{m}}=s^{(i)}_{m}$, the superscript $(i)$ denotes the component index. To enhance computational efficiency and reduce parametric complexity, we may constrain the metric tensor $\mathbf{M}_{m}$ to be a scalar matrix of the form $\mathbf{M}_{m}=s_{m}\mathbf{I}$, where $s_{m}>0$. This choice is geometrically equivalent to inducing a conformal deformation on the standard Euclidean space $\mathbb{R}^{D}$. In our geometric framework, this matrix $\mathbf{M}_{m}$ is not merely a scaling factor; it is interpreted as the metric tensor that defines the local geometry of the tangent space at position $m$. Its primary function is to modulate the inner product, thereby assigning a notion of semantic importance or saliency to the token at that position. The standard self-attention mechanism implicitly assumes a uniform Euclidean metric ($\mathbf{M}=\mathbf{I}$), treating all positions as equally important a priori. By introducing a position-dependent metric $\mathbf{M}_{m}$, we allow the model to differentiate the intrinsic significance of tokens. A larger scalar $s_{m}$ effectively amplifies the length of the query vector $\mathbf{q}_{m}$ and, consequently, the magnitude of its computed attention scores to all other keys. Thus, $s_{m}$ acts as a control knob for the token’s influence: a higher value signifies greater semantic importance, causing its query to have a stronger impact on the overall attention distribution.
While the general formulation $\mathbf{M}_{m}=s_{m}\mathbf{I}$ allows for arbitrary, position-specific importance, it does not explicitly encode a structural bias related to sequential order. To introduce such a prior, we can parameterize the scaling factor as an exponential function of the position index $m$. This is elegantly achieved by setting $\mathbf{M}_{m}=s^{m}\mathbf{I}$, where $s>0$ is a single, globally shared, learnable parameter. This highly parsimonious choice allows the model to learn a global decay rate ($s<1$) or a growth rate ($s>1$) of semantic importance, thereby systematically weighting tokens based on their absolute position in the sequence. We refer to this scheme as Conformal Exponential Scaling (CES). To impose the constraint $s<1$, we may proceed by defining it as $s=\frac{\exp(w)}{\exp(w)+\alpha}$, where $w$ is a learnable parameter and $\alpha$ denotes a hyperparameter. Alternatively, one might opt for $s=\exp(w)$, which permits $s>1$. For long sequences, we can decelerate the rate of exponential growth, by setting the parameter $s^{\log_{\beta}(m)}$. In addition, the learnable metric matrix of general form is given in the Appendix section.
@@ 2.3 A Principled Decomposition of High-Dimensional Rotations
2.3.1 Decomposition into Commuting $\mathbf{SO}(2)$ Subspaces
We can construct a basis for the Cartan subalgebra $\mathfrak{h}$ of $\mathfrak{so}(2k)$. This basis consists of $k$ matrices $\{H_{1},H_{2},\dots,H_{k}\}$:
$H_{1}=\begin{bmatrix}J&0&\cdots&0\\ 0&0&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&0\end{bmatrix},\quad H_{2}=\begin{bmatrix}0&0&\cdots&0\\ 0&J&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&0\end{bmatrix},\quad\dots\quad H_{k}=\begin{bmatrix}0&0&\cdots&0\\ 0&0&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&J\end{bmatrix}$,
where each $H_{i}$ is a $2k\times 2k$ block matrix, with all blocks being zero matrices except for the $i$-th diagonal block, which is the $2\times 2$ antisymmetric matrix $J$. The matrix $J$ is the generator of infinitesimal rotations in two dimensions:
@@ 2.3.1 Decomposition into Commuting $\mathbf{SO}(2)$ Subspaces
The Cartan subalgebra $\mathfrak{h}$ of $\mathfrak{so}(2k)$ is the linear span of this basis $\{H_{1},H_{2},\dots,H_{k}\}$. Any element $H$ within it can be expressed as:
@@
The exponential of $H$ is
@@
In fact, for any real antisymmetric matrix $X$, there exists an orthogonal matrix $Q$ such that $X=QHQ^{T}$, where $H$ is an element in the Cartan subalgebra $\mathfrak{h}$ of $\mathfrak{so}(2k)$, and thus $\exp(X)=Q\exp(H)Q^{T}$. Here, the rotation in the 2D subspace corresponds to setting $Q=I$, which simplifies the scenario but also limits its capacity for interactions across spatial dimensions. It is also worth noting that, under the aforementioned configuration, the proposed approach reduces to a conventional RoPE when setting the metric as $\mathbf{M}_{i}=\mathbf{I}$, thereby indicating that RoPE can be viewed as a special case of our framework.

### v3

@@ Abstract
The inherent permutation invariance of standard self-attention stems from its implicit assumption of a flat Euclidean geometry, where all spatial positions possess exchange symmetry. To fundamentally break this symmetry and endow the model with intrinsic geometric selectivity, we propose reframing attention as interactions on a curved Attention Manifold. In this framework, queries and keys are treated as vectors within the manifold’s intrinsic geometry, residing in the tangent bundle. Here, a Riemannian metric governs the local geometry, defining both the connection for parallel transport and the local inner product for similarity computation. In stark contrast, we propose that value vectors are sections of a separate, trivial vector bundle, decoupling the geometric alignment of $Q/K$ from the feature aggregation of $V$.
Our primary contribution is a novel, first-principles construction of this parallel transport using Lie group theory, where for 2D data, we derive a positional encoding from the exceptional isomorphism $\mathfrak{so}(4)\cong\mathfrak{su}(2)\oplus\mathfrak{su}(2)$. This allows us to factorize 4D rotations into two commuting isoclinic rotations aligned with the spatial dimensions. Furthermore, we introduce a Riemannian Attention Mechanism that strategically decouples metric scaling from the softmax normalization. Drawing inspiration from the physical principle that matter determines the geometry of space, we devise a dynamic metric scaling governed by the token embeddings themselves. This allows semantic content to locally modulate the geometry of the manifold, naturally inducing a content-aware focus that adaptively highlights salient regions. Experimental results demonstrate that our modules deliver significant performance improvements over strong baselines, and further evaluations on visual and large language models are underway.
@@ 1 Introduction
The transformer block operates on a sequence of tokens, each represented as a vector of fixed dimensionality $d_{\text{model}}$. For a sequence of $N$ tokens, the input to the block is a matrix $X\in\mathbb{R}^{N\times d_{\text{model}}}$, where each row corresponds to a token’s embedding. The transformer block processes this matrix through its core component, Multi-Head Self-Attention (MHSA). The MHSA mechanism is the cornerstone of the transformer, enabling it to aggregate information globally across all tokens in the input sequence. Unlike convolutional neural networks (CNNs), which rely on local receptive fields, or recurrent neural networks (RNNs), which process sequences sequentially, self-attention allows each token to attend to all other tokens simultaneously, capturing long-range dependencies efficiently. Specifically, the attention mechanism computes a score for each pair of tokens by taking the dot product of their query and key vectors, scaled to prevent large values that could destabilize training: $\text{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d_{k}}}\right)\mathbf{V}$. The dot product $\mathbf{QK}^{\top}\in\mathbb{R}^{N\times N}$ measures the similarity between the query vector of one token and the key vectors of all tokens, producing an attention score matrix. The softmax function normalizes the scores across the key dimension, converting them into weights that sum to $1$ for each query. This ensures that the attention mechanism focuses on the most relevant tokens. The resulting weights are used to compute a weighted sum of the value vectors, producing a new representation for each token that incorporates information from all other tokens. The output of the attention mechanism is a matrix, where each row is a context-aware representation of the corresponding input token. Moreover, the attention mechanism is permutation-equivariant, meaning the output does not depend on the order of the input tokens unless positional information is explicitly added.
@@
Motivation & Intuition. The inherent permutation invariance of standard self-attention stems from its implicit assumption of a flat Euclidean geometry. In such a homogeneous space, the metric structure is globally uniform ($\mathbf{M}=\mathbf{I}$), rendering all spatial positions geometrically indistinguishable and possessing a discrete translational symmetry. To fundamentally break this symmetry without relying on ad-hoc additive encodings, we propose embedding token sequences into a non-flat Riemannian Manifold.
In this curved geometry, the symmetry is naturally broken by the intrinsic structure of the manifold. Crucially, this Riemannian framework offers a dual advantage: it not only encodes spatial relationships via curvature but also quantifies the intrinsic importance of each token via the metric scale. By allowing the local metric to vary, we transform the "flat" attention landscape into a dynamic topography where salient features create "gravity wells", thereby endowing the self-attention mechanism with a geometrically principled selectivity.
@@ 2 Methodology
Consider a sequence comprising $N$ input tokens, with their embeddings represented as $\{\mathbf{x}_{i}\}_{i=1}^{N}$, where $\mathbf{x}_{i}\in\mathbb{R}^{D}$ denotes the $D$-dimensional embedding vector for the $i$-th token, devoid of positional information. In this paper, we adopt an alternative perspective: we posit the existence of a curved space $\mathcal{M}$ wherein the encoding of a token sequence corresponds to identifying a sequence of points $\{\mathbf{p}_{i}\}$, or a grid of points $\{\mathbf{p}_{ij}\}$ for image data, etc.
@@ 2.2 Parallel Transport
The above relationship of these matrices is what we must consider when predefining them. The Riemannian metric $\mathbf{M}_{i}$ is a symmetric positive matrix. To facilitate the model to effectively learn the linear transformations $\mathbf{T}_{i}$, we first decompose them into two components, including scaling and length-preserving in $\mathbb{R}^{D}$: $\mathbf{T}_{i}=\mathbf{S}_{i}\mathbf{R}_{i}$. The introduction of scaling part $\mathbf{S}_{i}$ is motivated by the characteristics of the curved space and the existence of Riemannian metric. The scaling part is represented by a symmetric positive matrix, and its simple form is a diagonal matrix or a scalar matrix $s\mathbf{I}$, where $\mathbf{I}$ is the identity matrix. For instance, we set $\mathbf{S}_{i}=\mathbf{M}_{i}^{-\frac{1}{2}}$, then we have $\mathbf{T}_{i}=\mathbf{M}_{i}^{-\frac{1}{2}}\mathbf{R}_{i}$. $\mathbf{R}_{i}$ is an orthogonal matrix used for length-preserving, and a rotation is naturally a candidate. A rotation in $D$-dimensional Euclidean space is a linear transformation that preserves lengths of vectors, and it can be represented by an $D\times D$ orthogonal matrix in the special orthogonal group $\mathbf{SO}(D)$.
@@
where $\langle\partial_{i},\partial_{i}\rangle_{\mathbf{p}_{m}}=s^{(i)}_{m}$, the superscript $(i)$ denotes the component index. To enhance computational efficiency and reduce parametric complexity, we may constrain the metric tensor $\mathbf{M}_{m}$ to be a scalar matrix of the form $\mathbf{M}_{m}=s_{m}\mathbf{I}$, where $s_{m}>0$. This choice is geometrically equivalent to inducing a conformal deformation on the standard Euclidean space $\mathbb{R}^{D}$. In our geometric framework, this matrix $\mathbf{M}_{m}$ is not merely a scaling factor; it is interpreted as the metric tensor that defines the local geometry of the tangent space at position $m$. Its primary function is to modulate the inner product, thereby assigning a notion of semantic importance or saliency to the token at that position.
2.3 Riemannian Attention Mechanism
The standard self-attention mechanism implicitly assumes a uniform Euclidean metric ($\mathbf{M}=\mathbf{I}$), treating all positions as equally important a priori. By introducing a position-dependent or embedding-dependent metric $\mathbf{M}_{m}$, we allow the model to differentiate the intrinsic significance of tokens. A larger scalar $s_{m}$ effectively amplifies the length of the query vector $\mathbf{q}_{m}$ and, consequently, the magnitude of its computed attention scores to all other keys. Thus, $s_{m}$ acts as a control knob for the token’s influence: a higher value signifies greater semantic importance, causing its query to have a stronger impact on the overall attention distribution.
Given the Riemannian metric tensor $\mathbf{M}_{m}$ at position $m$, the interaction between a query at $m$ and a key at $n$ is given by:
$\mathbf{q}_{m}^{\top}\mathbf{M}_{m}\mathbf{T}_{m}{\mathbf{T}_{n}}^{\!-1}\mathbf{k}_{n}=\langle\mathbf{T}_{m}^{-1}\mathbf{q}_{m},\;\mathbf{T}_{n}^{\!-1}\mathbf{k}_{n}\rangle.$
A direct incorporation of this term into the standard Transformer self-attention yields:
$\operatorname{Attn}=\text{softmax}\left(\frac{\mathbf{Q}^{\prime}\,\mathbf{K}^{\prime\top}}{\sqrt{d_{k}}}\right)\mathbf{V},$
where the $i$-th rows of $\mathbf{Q}^{\prime}$ and $\mathbf{K}^{\prime}$ are $\mathbf{T}_{i}^{-1}\mathbf{q}_{i}$ and $\mathbf{T}_{i}^{-1}\mathbf{k}_{i}$, respectively. Here, the transformation $\mathbf{T}_{i}=\mathbf{S}_{i}\mathbf{R}_{i}$ encapsulates both scaling and rotation. A primary motivation for introducing the scaling factor $s_{i}$ is to differentiate the intrinsic importance of different positions. However, the normalization property of the standard softmax function ($\sum p_{i}=1$) tends to counteract or "swallow" this scalar modulation, potentially diminishing its intended effect.
To address this, we propose the Riemannian Attention (RieAttn) mechanism. Instead of discarding softmax for unbounded alternatives like ReLU or Sigmoid, we retain the probabilistic interpretation of attention but decouple the metric scaling from the normalization step. We apply the rotational component $\mathbf{R}[\cdot]$ inside the softmax to determine directional alignment, while applying the metric scaling $\mathbf{S}$ outside the softmax as a gating term:
$\operatorname{RieAttn}=\left(\mathbf{S}\odot\text{softmax}\left(\frac{\mathbf{R}[\mathbf{Q}]\,\mathbf{R}[\mathbf{K}]^{\top}}{\sqrt{d_{k}}}\right)\right)\mathbf{V}=\left(\mathbf{L}\mathbf{L}^{\top}\odot\text{softmax}\left(\frac{\mathbf{R}[\mathbf{Q}]\,\mathbf{R}[\mathbf{K}]^{\top}}{\sqrt{d_{k}}}\right)\right)\mathbf{V}.$ (9)
Here, $\mathbf{R}[\cdot]$ denotes the application of the rotation $\mathbf{R}_{i}$ to the $i$-th row vectors. $\mathbf{L}=[s_{1},s_{2},\dots,s_{N}]^{\top}\in\mathbb{R}^{N}$ is the scale vector collecting scaling factors for all positions. Since we model the Riemannian metric as a conformal deformation, the scalar modulation for the interaction between any two positions naturally corresponds to the product of their respective scaling factors. Consequently, the global modulation matrix $\mathbf{S}$ emerges as the outer product $\mathbf{L}\mathbf{L}^{\top}$, ensuring a geometrically consistent and symmetric interaction structure.
We explore two distinct strategies for determining the scale vector $\mathbf{L}$:
1) Static Metric (Position-Dependent): The scaling factor is determined solely by the spatial location of the token. For 2D image data, this allows the model to learn a fixed "saliency map" of the image grid.
2) Dynamic Metric (Matter-Dependent): Motivated by the general relativistic principle that matter determines the geometry of spacetime, we propose that the metric at a specific location should be determined by its content (embedding) rather than just its coordinates.
For the dynamic setting, we model the scale vector as a function of the input embeddings: $\mathbf{L}=\mathfrak{g}(\text{Proj}(\mathbf{X}))$. The projection $\text{Proj}(\mathbf{X})=\mathbf{X}\mathbf{U}$, with learnable axes $\mathbf{U}\in\mathbb{R}^{D\times n_{h}}$ for each of the $n_{h}$ attention heads, effectively distills the content of each token into a lower-dimensional "mass" representation. The modulation function $\mathfrak{g}$ then maps this "mass" to a scaling factor that determines the local curvature of the Attention Manifold, implementing the principle that matter curves space. This mechanism offers significant interpretability: in NLP, it allows the manifold to contract around low-semantic function words, naturally dampening their influence; in Vision, it acts as an implicit saliency filter, amplifying the metric scale of foreground objects while suppressing the background. While $\mathfrak{g}$ can be standard activation functions like Sigmoid or ReLU, we empirically find the following formulation effective:
$\mathbf{L}=\mathfrak{g}(\text{Proj}(\mathbf{X}))=\sqrt{\alpha}\exp(\mathbf{X}\mathbf{U}),$ (10)
where $\alpha$ can be a learnable scalar and $\exp(\cdot)$ is applied element-wise. Consequently, for a single attention head, the scaling matrix $\mathbf{S}$ becomes:
[…]
where $t_{i}=\mathbf{X}_{i}\mathbf{U}$. Additionally, we can employ a more expressive Multilayer Perceptron (MLP) for the projection, defined as $\mathbf{L}=\mathfrak{g}\left(\text{MLP}(\mathbf{X})\right)$. For instance, $\mathbf{s}_{i}=\mathfrak{g}(\text{ReLU}(\mathbf{x}_{i}\mathbf{W}_{1})\mathbf{W}_{2})$, where $\mathbf{W}_{1}\in\mathbb{R}^{D\times d_{h}}$ and $\mathbf{W}_{2}\in\mathbb{R}^{d_{h}\times n_{h}}$, with $d_{h}$ representing the hidden dimension. This dynamic approach allows the attention mechanism to adaptively emphasize or suppress regions based on their semantic content, providing a robust inductive bias for semantic or visual understanding.
@@ 2.4 A Principled Decomposition of High-Dimensional Rotations
2.4.1 Decomposition into Commuting $\mathbf{SO}(2)$ Subgroups
The Lie algebra $\mathfrak{so}(2k)$ contains a subspace that is the direct sum of $k$ copies of $\mathfrak{so}(2)$:
$\mathfrak{h}=\bigoplus_{i=1}^{k}\mathfrak{so}(2)^{(i)}\subset\mathfrak{so}(2k).$
This subspace serves as the Cartan subalgebra $\mathfrak{h}$, for which we can construct a basis consisting of $k$ matrices $\{H_{1},H_{2},\dots,H_{k}\}$. Each basis matrix $H_{i}$ is built using the $2\times 2$ antisymmetric matrix $J$, which generates infinitesimal rotations in two dimensions:
@@ 2.4.1 Decomposition into Commuting $\mathbf{SO}(2)$ Subgroups
Each $H_{i}$ is a $2k\times 2k$ block-diagonal matrix, divided into $k$ blocks of size $2\times 2$. All blocks are zero except the $i$-th diagonal block, which is $J$:
$H_{1}=\begin{bmatrix}J&0&\cdots&0\\ 0&0&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&0\end{bmatrix},\quad H_{2}=\begin{bmatrix}0&0&\cdots&0\\ 0&J&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&0\end{bmatrix},\quad\dots,\quad H_{k}=\begin{bmatrix}0&0&\cdots&0\\ 0&0&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&J\end{bmatrix}$.
Any element $H$ within $\mathfrak{h}$ can be expressed as:
@@
From the perspective of Lie group theory, the exponential map $\exp(H)$ lies within the subgroup $SO(2)\times\cdots\times SO(2)\subset SO(2k)$. This subgroup consists of block-diagonal matrices, where each of the $k$ blocks is an element of $SO(2)$. Specifically, the exponential of $H$ is
@@
In fact, for any real antisymmetric matrix $X$, there exists an orthogonal matrix $Q$ such that $X=QHQ^{\top}$, where $H$ is an element in the Cartan subalgebra $\mathfrak{h}$ of $\mathfrak{so}(2k)$, and thus $\exp(X)=Q\exp(H)Q^{\top}$. Here, the rotation in the 2D subspace corresponds to setting $Q=I$, which simplifies the scenario but also limits its capacity for interactions across spatial dimensions. It is also worth noting that, under the aforementioned configuration, the proposed approach reduces to a conventional RoPE when setting the metric as $\mathbf{M}_{i}=\mathbf{I}$, thereby indicating that RoPE can be viewed as a special case of our framework.